OpenStudy (anonymous):

the area of the region enclosed by the graphs of y =x and y = x^2-3x+3 is? a) 2/3 b) 1 c) 4/3 d) 2 e) 14/3

OpenStudy (anonymous):

first write the equation of the parábola in vertex form

OpenStudy (anonymous):

ah should be y = (x-1)^2 + 1?

OpenStudy (anonymous):

i dont see why to put it in vertex form

OpenStudy (anonymous):

find intercept points: x^2-3x+3-x=0 x=1,x=3 \[\int\limits_{1}^{3}(x- x^2+3x-3)dx=4/3\] the answer c

OpenStudy (anonymous):

thanks! i got that far just mixed things up